9 Assume that you are considering the purchase of a 20-year, noncallable bond with an annual coupon rate of 9.5%. The bond has a face value of $1,000, and it makes semiannual interest payments. If you require an 10.7% nominal yield to maturity on this investment, what is the maximum price you should be willing to pay for the bond?
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

For this bond,

n = 20 * 2 = 40 semi-annual periods, C = 9.5% * $1000/2 = $47.5 (semi-annual), M = $1000, i = 10.7%/2 = 5.35% (semi-annual)

1- (1+0.0535)40 P = 47.5 * - 0.0535 1000 (1 + 0.053540

P = $777.45 + $124.34

P = $901.80